# Multi-Agent Reinforcement Learning for Connected and Automated Vehicles Control: Recent Advancements and Future Prospects

### v1

@@ IV-A Macro-micro energy optimization of CAVs using MARL
Efficient energy management in autonomous vehicles is crucial, as it not only ensures the optimal performance of each vehicle but also contributes to overall energy conservation [165]. This can be divided into two distinct yet complementary approaches: microscopic energy optimization, which concentrates on maximizing the efficiency of individual components within a single vehicle, and macroscopic energy optimization, which aims to enhance collective energy use across a fleet of autonomous vehicles.
@@
Microscopic energy optimization scrutinizes the performance of each component, aiming for optimal functioning. Traditional methods, which segregate powertrain, chassis control, and other controls, overlook the synergistic potential between these elements, potentially missing out on energy savings. To overcome this, Hua et al. present a collaborative strategy between two learning agents within the MADRL framework [166]. They conduct a sensitivity analysis to harmonize the settings for two learning agents, identifying an optimal mode that surpasses single-agent models by saving up to 4% more energy. Yang et al. develop a multi-objective energy management strategy for hybrid electric vehicles, leveraging MARL [167]. Their strategy simultaneously addresses fuel economy, battery charge maintenance, battery longevity, and ultracapacitor constraints, considering the engine–generator set (EGS) and hybrid energy storage system (HESS) as interactive agents. Shi et al. propose an energy management system (EMS) that uses independent Q-learning, a multi-agent reinforcement learning algorithm, to sustain battery charge levels and minimize hydrogen consumption, thereby optimizing energy for HESS[168]. Microscopic energy optimization enhances component efficiency but risks overlooking vehicle-wide dynamics. Macroscopic optimization addresses this by aiming to improve the energy efficiency across a fleet of autonomous vehicles.
@@
Macroscopic energy optimization has garnered increasing interest, particularly concerning traffic dynamics and their impact on energy efficiency. For instance, Wang et al. have developed a MARL-based strategy for hybrid electric vehicles that synchronizes powertrain management with car-following behaviors, thereby reducing energy use while maintaining safe distances [169]. Zhang et al. have designed a MARL method that simultaneously optimizes velocity and energy management for groups of HEVs, employing Markov games and LSTM networks for dynamic group understanding and an innovative asynchronous learning technique for inter-vehicle knowledge sharing[170]. Furthermore, Peng et al. utilize a MADDPG framework with two distinct agents for adaptive cruise control and energy management, respectively, achieving domain-specific optimization and enhancing the MADDPG’s effectiveness by more than 10% with an improved prioritized experience replay technique [171].
@@
Macro-micro energy optimization: The advancement of CAVs is set to enhance sustainable and efficient mobility, with macro-micro energy modeling playing a key role in precise energy consumption analysis, as depicted in Figure 10. This approach intertwines macro-level traffic analysis with micro-level hybrid electric vehicle (HEV) powertrain operations. By doing so, it allows for adaptive energy management that responds to real-time traffic conditions and powertrain status, optimizing the use of the electric motor and battery in various driving scenarios. Consequently, this integrated optimization strategy not only improves route efficiency and driving behavior but also leads to notable energy savings and cost reductions [172].
@@ IV-B Communication challenges
Current progress: Communication in MARL for CAVs is a critical element that underpins the efficiency and privacy of these intelligent transportation systems. The ability of CAVs to collaborate, share information, and make coordinated decisions hinges on effective communication protocols [173]. First, the development of advanced communication protocols tailored specifically for CAVs is a pressing need. Existing research focuses on protocols that enable efficient data sharing [174]. For instance, research endeavors have delved into the utilization of graph-based MARL algorithms [117], such as those explored in [130] and [132], to enable efficient and adaptive message exchange among agents. Second, the challenge of ensuring robust communication in dynamically evolving traffic scenarios represents another critical aspect that deserves meticulous attention [175]. For example, in [155], a variational auto-encoder algorithm combined with advanced multi-head attention mechanisms is proposed to extract and retain valuable information from neighboring agents while making efficient use of constrained communication resources. Last but not least, it is worth noting that managing unstable communication channels, particularly in terms of latency, emerges as a particularly intriguing consideration in the context of MARL for CAVs [96]. In [96], the double deep Q-learning algorithm is proposed to jointly train the agents to maximize the sum rate of V2N links while ensuring the desired packet delivery probability for each V2V link within specified latency constraints.
@@ IV-D Sim-to-real challenges
Deep MARL has achieved significant breakthroughs in various areas of the CAV domain. Given the constraints associated with real-world data collection, such as sample inefficiency and cost considerations, simulation environments have become invaluable for agent training (see Sec. III-D). Nevertheless, the disparity between the simulated and real-world settings can lead to a degradation in policy performance when models are transitioned to actual CAV scenarios [36]. In certain simulator platforms, a key objective is to enable the deployment of CAV agents with minimal disparities between their training data and real-world experiences [32, 11]. Simultaneously, other research initiatives are dedicated to enhancing safety, recognizing it as a primary obstacle to achieving real-world online training for complex self-driving car agents [35]. Conversely, it is crucial to explore advanced MARL techniques, including approaches like multi-agent offline RL [184] that leverage pre-existing offline data to update policies with online adaptations. Such future exploration can open new avenues for enhancing the generalization abilities and performance of CAVs in real-world scenarios.
@@ V Conclusion
In recent times, MARL has emerged as a focal point of interest within the realm of ITS. The capability of MARL to tackle intricate control and coordination challenges in CAVs has unlocked new horizons for the development of intelligent and efficient next-generation transportation systems. This review has undertaken a comprehensive exploration of the applications of MARL in the context of CAVs.
@@
    The review began with an overview of single-agent RL techniques and an extensive survey of the diverse landscape of MARL architectural variants.
@@
    Our analysis categorized these contributions based on different degrees of control, ranging from one-dimensional to two-dimensional and three-dimensional collaboration. Within this multifaceted arena, MARL has exhibited remarkable prowess, demonstrating its promising performance in various CAV control tasks. These include cooperative endeavors such as adaptive and cruise control, lane-changing maneuvers, on-ramp merging, traffic light coordination, and unscripted interactions, among others.
@@
    Amidst these accomplishments, we have also illuminated the formidable challenges that lie ahead. Designing and evaluating MARL methodologies within the CAV environment present their own set of complexities and intricacies. Balancing performance, scalability, and safety is no small feat.
This review stands as a valuable resource, poised to catalyze further research and the practical application of MARL within the domains of CAVs and ITS.

### v3

@@ IV-A Macro-micro energy optimization of CAVs using MARL
Efficient energy management in CAVs is crucial, as it not only ensures the optimal performance of each vehicle but also contributes to overall energy conservation [168, 169]. This can be divided into two distinct yet complementary approaches: microscopic energy optimization, which concentrates on maximizing the efficiency of individual components within a single vehicle, and macroscopic energy optimization, which aims to enhance collective energy use across a fleet of AVs.
@@
Microscopic energy optimization evaluates the performance of each component, aiming for optimal functioning. Traditional methods, which segregate powertrain, chassis control, and other controls, overlook the synergistic potential between these elements, potentially missing out on energy savings. To overcome this, Hua et al. present a collaborative strategy between two learning agents within the MADRL framework [170]. They conduct a sensitivity analysis to harmonize the settings for two learning agents, identifying an optimal mode that surpasses single-agent models by saving up to 4% more energy. Yang et al. develop a multi-objective energy management system (EMS) for hybrid electric vehicles (HEV) using a blend of game theory and RL within a MARL framework [171]. Their strategy simultaneously addresses fuel economy, battery charge maintenance, battery longevity, and ultracapacitor constraints, considering the engine–generator set (EGS) and hybrid energy storage system (HESS) as interactive agents. Shi et al. propose an EMS that uses IQL to sustain battery charge levels and minimize hydrogen consumption, thereby optimizing energy for HESS[172]. Microscopic energy optimization enhances component efficiency but risks overlooking vehicle-wide dynamics. Macroscopic optimization addresses this by aiming to improve the energy efficiency across a fleet of autonomous vehicles.
@@
Macroscopic energy optimization has garnered increasing interest, particularly concerning traffic dynamics and their impact on energy efficiency. For instance, Wang et al. have developed a MARL-based strategy for HEVs that synchronizes powertrain management with car-following behaviors, thereby reducing energy use while maintaining safe distances [173]. Zhang et al. have designed a MARL method that simultaneously optimizes velocity and energy management for groups of HEVs, employing Markov games and LSTM networks for dynamic group understanding and an innovative asynchronous learning technique for inter-vehicle knowledge sharing[174]. Furthermore, Peng et al. utilize a MADDPG framework with two distinct agents for adaptive cruise control and energy management, respectively, achieving domain-specific optimization and enhancing the MADDPG’s effectiveness by more than 10% with an improved prioritized experience replay [175].
@@
Macro-micro energy optimization: The advancement of CAVs is set to enhance sustainable and efficient mobility, with macro-micro energy modeling playing a key role in precise energy consumption analysis, as depicted in Fig. 10. This approach intertwines macro-level traffic analysis with micro-level HEV powertrain operations. By doing so, it allows for adaptive energy management that responds to real-time traffic conditions and powertrain status, optimizing the use of the electric motor and battery in various driving scenarios. Consequently, this integrated optimization strategy not only improves route efficiency and driving behavior but also leads to notable energy savings and cost reductions [176].
@@ IV-B Communication challenges
Current progress: Effective communication is pivotal in MARL for CAVs, enabling collaboration, information sharing, and coordinated decision-making through reliable protocols [177]. First, the development of advanced communication protocols tailored specifically for CAVs is a pressing need. Existing research focuses on protocols that enable efficient data sharing [178]. For instance, research endeavors have delved into the utilization of graph-based MARL algorithms [121], such as those explored in [134] and [136], to enable efficient and adaptive message exchange among agents. Second, the challenge of ensuring robust communication in dynamically evolving traffic scenarios represents another critical aspect that deserves meticulous attention [179]. For example, in [158], a variational auto-encoder algorithm combined with advanced multi-head attention mechanisms is proposed to extract and retain valuable information from neighboring agents while making efficient use of constrained communication resources. Last but not least, it is worth noting that managing unstable communication channels, particularly in terms of latency, emerges as a particularly intriguing consideration in the context of MARL for CAVs [96]. In [96], the double deep Q-learning algorithm is proposed to jointly train the agents to maximize the sum rate of V2N links while ensuring the desired packet delivery probability for each V2V link within specified latency constraints.
@@ IV-D Sim-to-real challenges
Deep MARL has achieved significant breakthroughs in various areas of the CAV domain. Given the constraints associated with real-world data collection, such as sample inefficiency and cost considerations, simulation environments have become invaluable for agent training (see Sec. III-D). Nevertheless, the disparity between the simulated and real-world settings can lead to a degradation in policy performance when models are transitioned to actual CAV scenarios [187]. In certain simulator platforms, a key objective is to enable the deployment of CAV agents with minimal disparities between their training data and real-world experiences [34]. Simultaneously, other research initiatives are dedicated to enhancing safety, recognizing it as a primary obstacle to achieving real-world online training for complex self-driving car agents [32]. Conversely, it is crucial to explore advanced MARL techniques, including approaches like multi-agent offline RL [188] that leverage pre-existing offline data to update policies with online adaptations. Such future exploration can open new avenues for enhancing the generalization abilities and performance of CAVs in real-world scenarios.
@@ V Conclusion
Recently, MARL has emerged as a focal point of interest within the realm of CAVs. The capability of MARL to tackle intricate control and coordination challenges in CAVs has unlocked new horizons for the development of intelligent and efficient next-generation transportation systems. This review has undertaken a comprehensive exploration of the applications of MARL in the control of CAVs.
@@
    The review began with an overview of single-agent RL techniques and an extensive survey of the diverse landscape of MARL architectural variants, facilitating a deeper understanding of applications to both individual and collective agent behaviors.
@@
    Our analysis categorized these contributions based on different degrees of control, ranging from one-dimensional to two-dimensional and three-dimensional collaboration. Within this multifaceted arena, MARL has exhibited remarkable prowess, demonstrating its promising performance in various CAV control tasks. These include cooperative endeavors such as platooning control, lane-changing maneuvers, on-ramp merging, traffic signal coordination, and unsignalized intersections, among others.
@@
    We highlighted significant challenges in designing and evaluating MARL methods within the CAV. Striking a balance between performance, scalability, and safety remains a complex challenge.
This review serves as a crucial resource, aimed at encouraging further research and the practical application of MARL within the field of CAVs.